OpenStudy (anonymous):

Write an equation of a line that is parallel to y = 3x – 5 and goes through the point (3, 1).

OpenStudy (anonymous):

Y=ax+b Parallel to Y=3x-5 === means a=3 Y=3x+b Goes through the point (3 ; 1) 1=3*3+b ==== b=1-9 =-8 Y=3x-8 done :)

OpenStudy (anonymous):

r u in connexus?